A particle vector is given by r=2ti+t^2j. Find: (a) r at t=1 and t=3, respectively;
(b) the displacement of particle from t=1 to t= 3.
(c) the path equation of the particle
1
Expert's answer
2020-05-13T10:14:11-0400

"\vec r=\vec r(t)=2t\vec i+t^2 \vec j=(2t,t^2)"


(a) "\vec r(1)=(2,1),\,\, \vec r(3)=(6,9)"


(b) "\Delta \vec r=\vec r(3)-\vec r(1)=(4,8)"


(c) "x=2t,\, y=t^2 \rightsquigarrow y=0.25x^2" is the path equation